# Resolvent-Based Optimisation for Invariant Flows

### v1

@@ Abstract
We present a robust optimisation framework for computing invariant solutions of wall-bounded flows by recasting the Navier–Stokes equations as a variational problem as established in Ashtari & Schneider, JFM (2023). The approach minimises the residual of the governing equations over a finite time horizon, seeking periodic or equilibrium solutions. A novel contribution is made by including a Galerkin projection onto a basis of divergence-free modes that satisfy the no-slip boundary conditions. This projection not only makes the variational framework applicable to wall-bounded flows but it also yields a low-order representation of the dynamics. The basis is derived from resolvent analysis, which provides an orthonormal set. We demonstrate the method on a 2D3C formulation of rotating plane Couette flow, obtaining exact equilibrium and periodic solutions consistent with direct numerical simulations. The conditioning of the optimisation problem is analysed in detail, showing that convergence rates depend on the stability properties of the targeted solutions. Finally, we highlight a direct link between the conditioning of the optimisation and the structure of the resolvent operator, suggesting a unifying perspective on both the efficiency of the optimisation and the dynamical significance of resolvent modes.
@@ 1 Introduction
For low-dimensional chaotic systems, the complex structure of strange attractors has been shown in Auerbach et al. (1987) to be determined by the set of Unstable Periodic Orbits (UPOs) embedded within the attractor, providing a “skeleton” for the dynamics. It was shown in Cvitanović (1988) and Cvitanović (1995) that building these UPOs up in the form of a weighted sum can yield the ergodic properties of the dynamics on the strange attractor (Artuso et al., 1990). This method is called cycle expansion theory. Analogously, exact nonlinear solutions to the Navier-Stokes equations have been numerically shown to exist, dubbed Exact Coherent Structures (ECSs) due to their close resemblance to coherent structures observed in experiment and Direct Numerical Simulation (DNS) studies. The exact nature of these ECSs are not yet fully understood, however, there is evidence that turbulent flows repeatedly shadow ECSs for a finite amount of time before moving away along one of the unstable manifolds of the solution towards a different ECSs (Suri et al., 2020; Krygier et al., 2021; Crowley et al., 2022). It was also demonstrated (Yalnız et al., 2021; Page et al., 2024) that using ECSs in a cycle expansion approach yields accurate statistics of the turbulent flow. The results do however generally rely on a tuning of the relative contribution of each ECS as opposed to deriving them exactly according cycle expansion theory (Wang et al., 2025). The numerical computation of ECSs was first shown in Nagata (1990) and Nagata (1997) where equilibrium and travelling wave solutions to the plane-Couette flow were computed. Since then there has been regular work to expand the sets of solutions for various flows (Waleffe, 1998, 2001; Faisst & Eckhardt, 2003; Wedin & Kerswell, 2004; Viswanath, 2007; Pringle & Kerswell, 2007; Gibson et al., 2008; Waleffe, 2009; Itano & Generalis, 2009; Wedin et al., 2009; Okino et al., 2010; Uhlmann et al., 2010; Willis et al., 2013; Nagata et al., 2021). A general review of the relevance of invariant solutions to fluid turbulence can be found in Kawahara et al. (2012).
@@
Global methods instead act on a full spatio-temporal field that already satisfies the required invariance, iteratively modifying it until the governing equations are solved to a prescribed tolerance. The Newton flow method introduced by Lan & Cvitanović (2004) formulates the problem through variational dynamics that converge exponentially to invariant solutions, with an over-relaxation factor mitigating the erratic behaviour of standard Newton iterations. Unlike shooting methods, these global approaches avoid exponential trajectory divergence and are generally more robust. However, the Newton flow method is not naturally matrix-free, as each iteration requires forming and solving an $N\times N$ Jacobian system, with $N$ the degrees of freedom. This makes the approach prohibitively costly for high-dimensional systems (Fazendeiro et al., 2010). Practical implementations must therefore employ GMRES, as in local methods. Despite these advantages, the method still suffers from sensitivity to initial guesses owing to the limited convergence radius of Newton iterations.
@@
An alternative global approach is to recast the search for invariant solutions as an optimisation problem rather than root finding. Here, the objective functional measures the total violation of the governing equations over the spatio-temporal domain. Introduced by Farazmand (2016) for equilibrium and travelling-wave solutions in 2D Kolmogorov flow, this method is formulated as the adjoint of the Newton descent of Lan & Cvitanović (2004). Crucially, it is naturally matrix-free and therefore well-suited to high-dimensional systems. The variational optimisation approach demonstrated markedly greater robustness to initial conditions, enabling convergence from a wider set of guesses and uncovering previously unknown solutions. More recently, Azimi et al. (2022), and Ashtari & Schneider (2023) extended the framework separately to periodic flows and to cases with no-slip boundary conditions, respectively. In the latter study, the Influence Matrix (IM) method (Kleiser & Schumann, 1980) was employed to enforce no-slip constraints consistently with the Pressure Poisson Equation (PPE), thereby avoiding boundary errors that arise when the PPE is solved without an explicit pressure condition.
@@
Despite its advantages, the variational optimisation method still faces challenges. While the IM method provides a clear strategy for equilibria in wall-bounded domains, its extension to temporally periodic fields is unclear, leaving open the problem of a general approach for periodic solutions in such settings. Moreover, the improved robustness compared with Newton-based methods comes at the cost of losing quadratic convergence. Convergence is initially rapid but slows considerably near the minimum, largely due to the use of gradient descent, which is notoriously inefficient in this regime (Nocedal & Wright, 2006). Algorithms incorporating curvature information could improve performance. Farazmand (2016) addressed this by adopting a hybrid strategy, initially using optimisation to approach the solution after which Newton–GMRES–hookstep is employed to accelerate convergence. More recently, Ashtari & Schneider (2023) improved efficiency of variational optimisation by employing an extrapolation technique with Dynamic Mode Decomposition (DMD). Nevertheless, the method remains hampered by a significantly slower convergence rate than Newton-based method, particularly as problem dimension increases. Hence, it is desirable to develop modifications that improve convergence without switching methods, while retaining compatibility with complex no-slip geometries.
@@
In this work, the variational optimisation framework is extended with a Galerkin projection, making it directly applicable to wall-bounded flows with no-slip boundaries. This projection enforces incompressibility and no-slip constraints while decoupling pressure and velocity. The basis is constructed from resolvent modes, which form a divergence-free, orthonormal set satisfying the boundary conditions. Our approach is similar to Li & Lasagna (2025), who used a projection-based optimisation for lid-driven cavity flow. In this work, however, resolvent modes replace SPOD modes, removing the need for fully resolved flow data and enabling the construction of exact solutions rather than reduced-order models. Resolvent analysis, a widely used tool for stability, control, and modelling (Gayme, 2010; Garnaud et al., 2013; Gómez et al., 2016; Beneddine et al., 2017; Symon et al., 2018; Gayme & Minnick, 2019; Jin et al., 2022), has been shown to provide dynamically significant bases for unstable solutions (Sharma et al., 2016) and to enable new invariant solutions via “projection-then-search” (Ahmed & Sharma, 2020). It has also been applied to reduced-order modelling, e.g. Taylor vortex flow in Barthel et al. (2021). Our work builds on these ideas by employing resolvent modes within a variational framework, where the Galerkin projection both enforces boundary constraints and yields a reduced-order model tailored to optimisation. Here, truncation is used not only to approximate turbulence, as in Barthel et al. (2021), but to enable computation of exact invariant solutions and accelerate convergence. To further address the slow convergence of variational optimisation, we depart from gradient-descent schemes (Farazmand, 2016; Azimi et al., 2022; Ashtari & Schneider, 2023) and instead use quasi-Newton algorithms such as L-BFGS and conjugate gradient, which achieve substantially faster convergence.
@@
The remainder of the paper is organised as follows. Section 2 introduces the variational optimisation methodology and the Galerkin projection used to enforce incompressibility and no-slip boundary conditions. Section 3 presents the derivation of the resolvent modes forming the projection basis. Section 4 introduces the 2D3C formulation of Rotating Plane Couette Flow (RPCF) and features a basic analysis of the behaviour of the flow at various Reynolds number regimes. Section 5 then demonstrates the projected variational optimisation methodology using resolvent analysis, featuring equilibrium and periodic solutions. The underlying mechanics governing the conditioning of the optimisation is then discussed in section 6, providing a novel link between resolvent analysis and the convergence rate. Finally, section 7 concludes with a discussion of the main results of this work and the possible avenues for future work.
@@ 2 Methodology
In section 2.1, the variational optimisation methodology for wall-bounded flows is described. The methodology is applicable to general wall-bounded flows, however here it is specialised for flows with a single inhomogeneous direction. This is to be more in line with the derivation in Ashtari & Schneider (2023). In section 2.2 the Galerkin projection is introduced as a way to solve to issues that arise due to the presence of no-slip boundary conditions. A more verbose derivation and discussion of the implications can be found in Burton (2025).
2.1 Variational Optimisation for Wall-Bounded Periodic Flows
@@
where $\mathit{Re}$ denotes the Reynolds number. The exact definition is delays until section 4 where the particular flow studied in detail in this work is defined. The velocity $\bm{u}$ is an $3$-dimensional vector field defined over a compact $3$-dimensional space, denoted as $\Omega=[0,\,L_{x}]\times[-1,\,1]\times[0,\,L_{z}]$ with a boundary denoted as $\partial\Omega$. $p$ denotes the scalar pressure field defined on the same space as $\bm{u}$. The coordinates for $\Omega$ are the single inhomogeneous wall-normal direction ($y$) normalised with respect to the channel height, and the streamwise and spanwise directions ($x$ and $z$ respectively) which are statistically homogeneous. Using these Cartesian coordinates the Laplace operator is defined as $\bm{\Delta}=\partial^{2}/\partial x^{2}+\partial^{2}/\partial y^{2}+\partial^{2}/\partial z^{2}$. Homogeneity of the flow is modelled with periodic boundary conditions in the streamwise and spanwise directions, while the no-slip boundary conditions are imposed at the wall
@@
where $\bm{u}_{w}$ is velocity of the wall. The velocity field $\bm{u}$ is also defined over a finite time horizon fixed by the orbit period $T$, thus $t\in\left[0,\,T\right)$. The combination of the spacial and temporal domains is defined as $\Omega_{t}=\Omega\crossproduct\left[0,\,T\right)$. The inner-product on this space is defined as
@@
The elements of $\mathcal{P}_{T}$ are state-space loops that do not necessarily obey (1a). In general $T$ is not known a priori and should be included as part of the optimisation problem. The goal is to derive a method that converges to the subset of $\mathcal{P}_{T}$ that obey (1a). This can be done by introducing a mapping $\mathcal{R}\colon\mathcal{P}_{T}\to\mathbb{R}_{\geq 0}$ such that $\mathcal{R}\left[\bm{u}\right]=0$ if and only if $\bm{u}\in\mathcal{P}_{T}$ and $\bm{u}$ satisfies (1a), otherwise $\mathcal{R}>0$. This is equivalent to defining an objective function whose global minima ($\mathcal{R}=0$) correspond to periodic solutions of (1a) that are divergence-free and obey the no-slip and periodicity boundary conditions, called here the global residual. To obtain the global residual, first a local residual that quantifies the violation of the Navier-Stokes equations at every point in $\Omega_{t}$ needs to be defined as follows
@@
where the functional $\mathcal{R}\left[\bm{u}\right]$ is called the global residual. Geometrically, minimising $\mathcal{R}$ is equivalent to continuously deforming a loop in the state-space of the flow field, constrained to the linear subspace that define the boundary conditions and incompressibility conditions. This process terminates when the rate of change of the state vector and the right-hand side of (1a) are as closely aligned as possible.
@@
as well as $\bm{r}$ obeying the periodic boundary conditions. The details of the derivation for (8) and the source of the extra constraints in (9) are given in Appendix A.
@@
To implement a gradient-based optimisation of (6) it is also necessary to have a gradient of $T$. In this work, however, the variable $\omega=2\pi/T$ is preferred. The variable $\omega$ is called the fundamental frequency and represents the lowest frequency oscillation that is permissible in the finite time window defined by the period $T$. The gradient of $\mathcal{R}$ with respect to $\omega$ is given by
of which the detailed derivation is given in Appendix A. Comparing (8) and (9) against (1) and (2), it can be seen that the optimisation can be viewed as a new set of “variational dynamics”, some details of which are discussed in appendix B.

### v2

@@ Abstract
We present a robust optimisation framework for computing invariant solutions of wall-bounded flows by recasting the Navier–Stokes equations as a variational problem as established in Ashtari & Schneider, JFM (2023). The approach minimises the residual of the governing equations over a finite time horizon, seeking periodic or equilibrium solutions. A novel contribution is made by including a Galerkin projection onto a basis of divergence-free modes that satisfy the no-slip boundary conditions. This projection not only makes the variational framework applicable to wall-bounded flows but it also yields a low-order representation of the dynamics. The basis is derived from resolvent analysis, which provides an orthonormal set. We demonstrate the method on a streamwise invariant formulation of rotating plane Couette flow, obtaining exact equilibrium and periodic solutions consistent with direct numerical simulations. The conditioning of the optimisation problem is analysed in detail, showing that convergence rates depend on the stability properties of the targeted solutions. Finally, we highlight a direct link between the conditioning of the optimisation and the structure of the resolvent operator, suggesting a unifying perspective on both the efficiency of the optimisation and the dynamical significance of resolvent modes.
@@ 1 Introduction
For low-dimensional chaotic systems, the complex structure of strange attractors has been shown in Auerbach et al. (1987) to be determined by the set of Unstable Periodic Orbits (UPOs) embedded within the attractor, providing a “skeleton” for the dynamics. It was shown in Cvitanović (1988) and Cvitanović (1995) that building these UPOs up in the form of a weighted sum can yield the ergodic properties of the dynamics on the strange attractor (Artuso et al., 1990). This method is called cycle expansion theory. Analogously, exact nonlinear solutions to the Navier-Stokes equations have been numerically shown to exist, dubbed Exact Coherent Structures (ECSs) due to their close resemblance to coherent structures observed in experiment and Direct Numerical Simulation (DNS) studies. The exact nature of these ECSs is not yet fully understood, however, there is evidence that turbulent flows repeatedly shadow ECSs for a finite amount of time before moving away along one of the unstable manifolds of the solution towards a different ECSs (Suri et al., 2020; Krygier et al., 2021; Crowley et al., 2022). It was also demonstrated (Yalnız et al., 2021; Page et al., 2024) that using ECSs in a cycle expansion approach yields accurate statistics of the turbulent flow. The results do however generally rely on a tuning of the relative contribution of each ECS as opposed to deriving them exactly according cycle expansion theory (Wang et al., 2025). The numerical computation of ECSs was first shown in Nagata (1990) and Nagata (1997) where equilibrium and travelling wave solutions to the plane-Couette flow were computed. Since then there has been regular work to expand the sets of solutions for various flows (Waleffe, 1998, 2001; Faisst & Eckhardt, 2003; Wedin & Kerswell, 2004; Viswanath, 2007; Pringle & Kerswell, 2007; Gibson et al., 2008; Waleffe, 2009; Itano & Generalis, 2009; Wedin et al., 2009; Okino et al., 2010; Uhlmann et al., 2010; Willis et al., 2013; Nagata et al., 2021). A general review of the relevance of invariant solutions to fluid turbulence can be found in Kawahara et al. (2012).
@@
Global methods instead act on a full spatio-temporal field that already satisfies the periodicity constraint, iteratively modifying it until the governing equations are satisfied to within a prescribed tolerance. The Newton flow method introduced by Lan & Cvitanović (2004) formulates the problem through variational dynamics where the solution residual decays exponentially (in the limit of continuous deformation) to invariant solutions, with an over-relaxation factor mitigating the erratic behaviour of standard Newton iterations. Unlike shooting methods, these global approaches avoid exponential trajectory divergence and are generally more robust. However, the Newton flow method is not naturally matrix-free, as each iteration requires forming and solving an $N\times N$ Jacobian system, with $N$ the degrees of freedom for both space and time. This makes the approach prohibitively costly for high-dimensional systems (Fazendeiro et al., 2010). Practical implementations must therefore employ GMRES, as in local methods. Despite these advantages, the method still suffers from sensitivity to initial guesses owing to the limited convergence radius of Newton iterations.
Fourier–Galerkin (or harmonic-balance) formulations offer an efficient global approach for computing time-periodic invariant solutions without long integrations. By projecting the Navier–Stokes equations onto a truncated Fourier basis in time, the problem reduces to a nonlinear algebraic system for the harmonic coefficients. Recent work has demonstrated the effectiveness of this strategy for periodic and quasi-periodic flows (Sierra-Ausin et al., 2022), its extension to adjoint-based sensitivity and stability analysis (Sierra et al., 2021), and its application to nonlinear input–output analysis capturing triadic energy transfers (Rigas et al., 2021). These Fourier–Galerkin approaches provide a complementary route to invariant solutions, trading time-marching for a frequency-domain nonlinear solve while naturally incorporating harmonic interactions and sensitivities.
@@
An alternative global approach is to recast the search for invariant solutions as an optimisation problem rather than root finding. Here, the objective functional measures the total violation of the governing equations over the spatio-temporal domain. Introduced by Farazmand (2016) for equilibrium and travelling-wave solutions in 2D Kolmogorov flow, this method is formulated as the adjoint of the Newton descent of Lan & Cvitanović (2004). Crucially, it is naturally matrix-free and therefore well-suited to high-dimensional systems. The variational optimisation approach demonstrated markedly greater robustness to initial conditions, enabling convergence from a wider set of guesses and uncovering previously unknown solutions. More recently, Azimi et al. (2022) extended the framework to periodic flows, and Ashtari & Schneider (2023) applied it to finding equilibrium in wall-bounded problems by enforcing no-slip constraints through the Influence Matrix (IM) method (Kleiser & Schumann, 1980), which maintains consistency with the Pressure Poisson Equation (PPE) and avoids boundary errors arising from the lack of an explicit pressure condition.
@@
Despite its advantages, the variational optimisation method still faces challenges. While the IM method provides a clear strategy for equilibria in wall-bounded domains, its extension to temporally periodic fields has not yet been published in the literature, leaving open the problem of a general approach for periodic solutions in such settings. Moreover, the improved robustness compared with Newton-based methods comes at the cost of losing quadratic convergence. Convergence is initially rapid but slows considerably near the minimum, largely due to the use of gradient descent, which is notoriously inefficient in this regime (Nocedal & Wright, 2006). Algorithms incorporating curvature information could improve performance. Farazmand (2016) addressed this by adopting a hybrid strategy, initially using optimisation to approach the solution after which Newton–GMRES–hookstep is employed to accelerate convergence. More recently, Ashtari & Schneider (2023) improved efficiency of variational optimisation by employing an extrapolation technique with Dynamic Mode Decomposition (DMD). Nevertheless, the method remains hampered by a significantly slower convergence rate than Newton-based methods, particularly as problem dimension increases. Hence, it is desirable to develop modifications that improve convergence without switching methods, while retaining compatibility with complex no-slip geometries.
@@
In this work, the variational optimisation framework is extended with a Galerkin projection, providing a general methodology for wall-bounded periodic flows with no-slip boundaries. The form of this problem closely resembles the general harmonic balance method described in Sierra-Ausin et al. (2022). This projection enforces incompressibility and no-slip constraints while decoupling pressure and velocity. The basis is constructed with resolvent modes, which form a divergence-free orthonormal set satisfying the boundary conditions. Our approach is similar to Li & Lasagna (2025), who used a projection-based optimisation for lid-driven cavity flow. In this work, however, resolvent modes replace SPOD modes, removing the need for fully resolved flow data and enabling the construction of exact solutions rather than reduced-order models. Resolvent analysis, a widely used tool for stability, control, and modelling (Gayme, 2010; Garnaud et al., 2013; Gómez et al., 2016; Beneddine et al., 2017; Symon et al., 2018; Gayme & Minnick, 2019; Jin et al., 2022), has been shown to provide dynamically significant bases for unstable solutions (Sharma et al., 2016; Rosenberg & McKeon, 2019) and to enable new invariant solutions via “projection-then-search” (Ahmed & Sharma, 2020). It has also been applied to reduced-order modelling, e.g. Taylor vortex flow in Rosenberg & Mckeon (2019) and Barthel et al. (2021). Our work builds on these ideas by employing resolvent modes within a variational framework, where the Galerkin projection both enforces boundary constraints and effectively pre-conditions the optimisation. In this setting, modal truncation is used not merely to approximate turbulence, as in Barthel et al. (2021), but to enable computation of exact invariant solutions while simultaneously improving the conditioning of the optimisation and thereby accelerating convergence. To further address the slow convergence inherent to variational optimisation, we utilise two strategies. First, we replace the gradient-descent schemes of Farazmand (2016); Azimi et al. (2022); Ashtari & Schneider (2023) with quasi-Newton algorithms such as L-BFGS and conjugate gradient, which incorporate curvature information. We demonstrate a substantially faster convergence rate near the minimum compared to gradient descent. Second, we exploit the Galerkin–resolvent framework itself by controlling the truncation of the resolvent basis we directly influence the conditioning of the optimisation problem, linking the achievable convergence rate to the singular values and modes of the resolvent operator. This dual use of the projection framework thus addresses both the difficulty of enforcing no-slip constraints in wall-bounded flows and the slow convergence that has limited the practical utility of variational optimisation.
@@
The remainder of the paper is organised as follows. Section 2 introduces the variational optimisation methodology and the Galerkin projection used to enforce incompressibility and no-slip boundary conditions. Section 3 presents the derivation of the resolvent modes forming the projection basis. Section 4 introduces the 2D3C formulation of Rotating Plane Couette Flow (RPCF) and features a basic analysis of the behaviour of the flow at various Reynolds number regimes. Section 5 then demonstrates the projected variational optimisation methodology using resolvent analysis, featuring equilibrium and periodic solutions. The underlying mechanics governing the conditioning of the optimisation is then discussed in section 6, providing a novel link between resolvent analysis and the convergence rate related to the formulation discussed in Mons & Marquet (2021). Finally, section 7 concludes with a discussion of the main results of this work and the possible avenues for future work.
@@ 2 Methodology
In section 2.1, the variational optimisation methodology for wall-bounded flows is described. The methodology is specialised for planar wall-bounded flows such as channel flow or Couette flow. This is to be more in line with the derivation in Ashtari & Schneider (2023). In section 2.2 the Galerkin projection is introduced as a way to solve to issues that arise due to the presence of no-slip boundary conditions. A more detailed derivation and discussion of the implications can be found in Burton (2025).
2.1 Variational Optimisation
@@
where $\mathit{Re}$ denotes the Reynolds number of the particular flow. The coordinates for the domain, denoted as $\Omega$, are the single inhomogeneous wall-normal direction ($y$) normalised with respect to the half channel height denoted as $h$, and the streamwise and spanwise directions ($x$ and $z$, respectively) which are statistically homogeneous. The domain is thus defined as $\Omega=[0,\,L_{x}]\times[-1,\,1]\times[0,\,L_{z}]$ with $L_{x}$ and $L_{z}$ denoting the streamwise and spanwise length of the domain, respectively, normalised with respect to the channel height. The velocity $\bm{u}$ is a $3$-dimensional vector field defined over the $3$-dimensional space $\Omega$, and the pressure field, denoted with $p$, is a scalar field over the same domain. The Laplace operator of functions in this domain is defined as $\bm{\Delta}=\partial^{2}/\partial x^{2}+\partial^{2}/\partial y^{2}+\partial^{2}/\partial z^{2}$. Homogeneity is modelled with periodic boundary conditions in the streamwise and spanwise directions, while the no-slip boundary conditions are imposed at the wall
@@
where $\bm{u}_{w}$ is velocity of the wall. A non-zero value for $\bm{u}_{w}$ can be used to represent a moving wall such as in Couette flow, and zero when the flow is driven through some other mechanism such as a pressure gradient in channel flow. The velocity field $\bm{u}$ is also defined over a finite time horizon fixed by the orbit period $T$, thus $t\in\left[0,\,T\right)$. The combination of the spatial and temporal domains is defined as $\Omega_{t}=\Omega\crossproduct\left[0,\,T\right)$. The inner-product on this space is defined as
@@
The elements of $\mathcal{P}_{T}$ are state-space loops that do not necessarily obey (1a). In general $T$ is not known a priori and should be included as part of the search for a solution. The goal is to derive a method that converges to elements of the subset of $\mathcal{P}_{T}$ that obey (1a). This can be done by introducing a mapping $\mathcal{R}\colon\mathcal{P}_{T}\to\mathbb{R}_{\geq 0}$ such that $\mathcal{R}\left[\bm{u}\right]=0$ if and only if $\bm{u}\in\mathcal{P}_{T}$ and $\bm{u}$ satisfies (1a); otherwise, $\mathcal{R}>0$. This is equivalent to defining an objective function whose global minima ($\mathcal{R}=0$) correspond to periodic solutions of (1a) that are divergence-free and obey the no-slip and periodicity boundary conditions, called here the global residual. To obtain the global residual, first a local residual that quantifies the violation of the Navier-Stokes equations at every point in $\Omega_{t}$ needs to be defined as follows
@@
Geometrically, minimising $\mathcal{R}$ is equivalent to continuously deforming a state-space loop as depicted in figure 1 such that the vectors $\partialderivative*{\bm{u}}{t}$ and $\gradient{p}+\left(\bm{u}\cdot\gradient{\bm{u}}\right)-\frac{1}{\mathit{Re}}\bm{\Delta}\bm{u}$ align as closely as possible, while being constrained to the linear subspace that define the boundary conditions and incompressibility conditions. This process terminates when the rate of change of the state vector and the right-hand side of (1a) are as closely aligned as possible.
@@
as well as $\bm{r}$ obeying the periodic boundary conditions. The details of the derivation for (8) and the source of the extra constraints in (9) are given in appendix A.1.
To implement a gradient-based optimisation of (6) it is also necessary to have a gradient of $\mathcal{R}$ with respect to $T$. In this work, however, the fundamental frequency, defined as $\omega=2\pi/T$, is preferred. The fundamental frequency represents the lowest frequency oscillation that is permissible in the finite time window defined by the period $T$. The gradient of $\mathcal{R}$ with respect to $\omega$ is given by
of which the detailed derivation is given in appendix A.2.